OpenStudy (tanabugg):

Write the equation in standard form using integers. y= -2/3x - 5

OpenStudy (steve816):

First of all, standard form is Ax + By = C

OpenStudy (steve816):

Rearrange the equation by adding -2/3x to both sides. You would get\[\large \frac{ 2 }{ 3 }x +y = -5\]

OpenStudy (steve816):

Now, in order to get the equation into integers, multiply by 3 to ALL terms, and that will cancel the fraction out :)

OpenStudy (tanabugg):

Wait I don't get the last part ?? @steve816

OpenStudy (steve816):

SImply multiply everything by 3!

OpenStudy (tanabugg):

okay so 2/3 * 3 = 2 and -5 * 3= -15 is that it?? @steve816

OpenStudy (tanabugg):

and it'd look like 2x + y = -15 ??

OpenStudy (tanabugg):

@triciaal

OpenStudy (steve816):

It would be 2x + 3y = -15 ^^ that is your answer!